It is known that 65% of Empire plan customers order their prescriptions via mail-order at least once a year. Suppose we randomly select 6 customers.

What is the probability that at most one has mail-ordered in the last year?
Answer #1

Solution

Given that ,

p = 0.65

n = 6

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X 1) = P(X = 0) + P(X = 1)

= (6 C 0 ) * 0.650 * (0.35)0  + (6 C 1) * 0.651 * (0.35)5

= 0.0223

Probability = 0.0223